# RealityCraft: An In-Situ CAD+CAM Interface for Novices via Scene-Aware Augmented Reality

### v1

@@ 6.2.3. Surveys
Results of the UEQ and NASA-TLX surveys.
We obtained an astonishing score of for the standard SUS survey (Figure 12). This score falls within the Grade B range, which suggests that the tool is perceived as highly usable, intuitive, and easy to navigate, with minimal friction during interactions. The UEQ results (Figure 13-left) also showed strong positive evaluations across both pragmatic (usability and task-related aspects) and hedonic (appeal and enjoyment) qualities. Specifically, the system scored 1.93 on pragmatic quality and 2.07 on hedonic quality, placing RealityCraft in the top 10% of UEQ scores (Laugwitz et al., 2008). A small portion of participants (shown in red) rated the system as confusing, which suggests that certain aspects of the interface and onboarding may require further refinement to enhance clarity, particularly for new users. Nevertheless, participants responded positively to most elements, particularly its supportiveness and ease of use.
As shown on the right side of Figure 13, NASA-TLX scores revealed a low overall cognitive workload, with most participants reporting minimal mental and physical effort. Specifically, participants also found the effort and frustration levels to be lower than average, while performance satisfaction was rated highly.
@@ 6.3. Qualitative Feedback
Open-ended questions in our survey, such as ”What did you like most about using the tool?”, ”What challenges did you face while using the tool?”, and ”What aspects of the tool would you most like to see improved?”, allowed us to gather deeper insights into the system’s strengths and areas for improvement.
Intuitiveness and Efficiency. Participants praised the system’s real-time AR design capabilities, highlighting the ability to visualize objects in situ. For instance, one participant noted the benefit of ”visualizing the objects I designed in my real-world work environment” (P4). Another emphasized the ease of creating designs with hand gestures, stating, ”the simplicity of creating an object from imagination with my own hands” (P3). The self-explanatory interface and the use of primitive geometries were also well received, with one user describing it as ”easy to use, rich but also straightforward” (P2) and another appreciating the ability to ”model with simple geometries” (P7). One participant commented on the integrated 3D printer support: ”Magical that designing and printing is in a unified package. I can print designs easily” (P6).
Comparative Advantage. Compared to traditional CAD tools, RealityCraft’s AR-based, hands-on design environment offered unique advantages. Users appreciated the ability to design in real-time while interacting with the physical world. For example, one participant praised the process of ”designing objects in the room” (P5), while another highlighted the benefit of being able ”to see all views of an object” (P1). The use of hand gestures for modeling was another frequently mentioned advantage.
Areas for Improvement. Several participants pointed out limitations with the precision of hand tracking, particularly during fine-tuned operations. This issue was more noticeable in free movement mode compared to snap-to-grid mode. Some participants suggested the addition of more precise object manipulation methods, such as ”numerically adjusting their positions” (P6). Additionally, minor bugs were reported, including issues with depth occlusion, as one participant observed, ”Objects sometimes cannot be seen due to z-fighting issues in depth occlusion mode” (P4). Others pointed to occasional interface challenges, such as ”UI issues, e.g., unwanted clicks, and not very precise hand/finger tracking” (P2).
@@ 7. Discussion
In this section, we evaluate RealityCraft’s effectiveness and future potential. We consider its suitability for both novice and expert users and discuss limitations related to handling complex 3D geometries. We also address the impact of the AR form factor on user experience and the limitations of hand gesture recognition. Additionally, we anticipate future advancements that could improve the system’s capabilities and user interaction.
@@
Our tool was designed with novices in mind, providing an intuitive interface that simplifies both the design and fabrication of models. However, we envision that RealityCraft can also be beneficial to experts. Practitioners and professional designers can use it for rapid prototyping, concept development, and interactive presentations. The ease of use and flexibility of our system can streamline workflows, allowing experts to focus on more complex design tasks without being bogged down by technical logistics. In the future, we believe that our system can be adopted by educational institutions for teaching design principles, by industries for quick mock-ups or collaborative sessions, and by hobbyists seeking to explore 3D modeling and fabrication.
@@
In Section 5, we demonstrated our tool’s usefulness through various design examples, each showing a different use case. These objects featured distinct geometries, ranging from small details to larger-scale structures. For instance, the candy bowl model in Figure 10d is made of 32 curved lines, each made of 8 cuboids, adding up to a total of 256 basic building blocks (if none of the blocks are combined) and 2048 vertices. Our system supports up to 300 basic building blocks and a total number of vertices of 20000 in real-time. However, we observed a noticeable lag when exceeding these numbers, especially in combination operations. We anticipate this limitation can be addressed with future headsets offering higher processing power. Alternatively, more powerful headsets, e.g., Varjo XR-4121212https://varjo.com/products/xr-4/, Accessed: 4.08.2024, which is connected to a desktop computer, can support compute-intensive tasks.
While the system can create a variety of complex geometries, certain tasks remain challenging due to the lack of more advanced tools such as extrusion of 3D geometries from 2D blueprints (e.g., sketched shapes, text), parametric curves, reference plane tools, lofting, filleting, object-to-object snap-to-surface functionality, freeform design tools similar to those in Autodesk 3ds Max131313https://www.autodesk.com/products/3ds-max, Accessed: 4.08.2024, or a collaborative structure allowing multiple users to work on the same project similar to Gravity Sketch or adding physical engagement with local users and their surroundings, similar to HoloBots by Ihara et al. (Ihara et al., 2023). These features were not implemented as our system targets novices and aims for simplicity.
@@
We believe that in addition to 3D printing, RealityCraft could be adapted for other fabrication methods such as molding, CNC machining, or laser cutting in the future. The primary difference lies in converting designs to machine-specific instructions rather than generating a .gcode file for a 3D printer. While laser cutting typically involves 2D plates, we can leverage the work by Roumen et al., particularly (Baudisch et al., 2019), which enables users to design 3D geometries using 2D plates.
@@
MR headsets can be bulky; however, in our user study, participants did not report major problems due to the form factor. We used the Quest 3 because it is lightweight, making it easy to use. In the future, we imagine AR technology becoming more ubiquitous while offering a more streamlined and integrated experience with the advent of AR glasses. Future devices are likely to be more compact, resembling ordinary eyewear, thereby enhancing user comfort and accessibility. These advancements could lead to broader adoption and more immersive systems.
@@
While we are constrained by Meta’s hand gesture APIs, which are known to improve over time141414https://www.techradar.com/computing/virtual-reality-augmented-reality/your-meta-quest-3-is-getting-a-hand-tracking-upgrade-that-could-unlock-foot-tracking, Accessed: 5.08.2024, there are still issues such as improper functionality in dimly lit areas or limited precision of tracking compared to controllers or the system being exclusive to Meta headsets.
@@ 8. Conclusion
In this paper, we presented RealityCraft, a new 3D modeling and fabrication tool through AR, designed for novices. Our approach lets users design artifacts with a simple primitive geometry manipulation design logic, with intuitive hand gesture controls, on the intended space of the artifact, and with seamless integration for the fabrication process. We showed the various features present in RealityCraft from selecting the workspace to design and fabrication features. Using these features, we presented ten creations for different scenarios, each varying in complexity and practical applications, hinting at what can be done with RealityCraft. The user study we conducted revealed that the system was highly effective for novices. We hope that our project and its open-sourced codebase will stimulate further innovation and exploration within the communities engaged with HCI, design, and fabrication.
@@ Acknowledgements.
- Cao et al. (2023) Jacky Cao, Kit-Yung Lam, Lik-Hang Lee, Xiaoli Liu, Pan Hui, and Xiang Su. 2023. Mobile Augmented Reality: User Interfaces, Frameworks, and Intelligence. ACM Comput. Surv. 55, 9 (Jan. 2023), 189:1–189:36. https://doi.org/10.1145/3557999
@@
- Chidambaram et al. (2021) Subramanian Chidambaram, Hank Huang, Fengming He, Xun Qian, Ana M Villanueva, Thomas S Redick, Wolfgang Stuerzlinger, and Karthik Ramani. 2021. ProcessAR: An augmented reality-based tool to create in-situ procedural 2D/3D AR Instructions. In Proceedings of the 2021 ACM Designing Interactive Systems Conference (DIS ’21). Association for Computing Machinery, New York, NY, USA, 234–249. https://doi.org/10.1145/3461778.3462126
- Dogan et al. (2022) Mustafa Doga Dogan, Patrick Baudisch, Hrvoje Benko, Michael Nebeling, Huaishu Peng, Valkyrie Savage, and Stefanie Mueller. 2022. Fabricate It or Render It? Digital Fabrication vs. Virtual Reality for Creating Objects Instantly. In Extended Abstracts of the 2022 CHI Conference on Human Factors in Computing Systems (New York, NY, USA, 2022-04-28) (CHI EA ’22). Association for Computing Machinery, 1–4. https://doi.org/10.1145/3491101.3516510
- Dogan et al. (2024) Mustafa Doga Dogan, Eric J. Gonzalez, Karan Ahuja, Ruofei Du, Andrea Colaço, Johnny Lee, Mar Gonzalez-Franco, and David Kim. 2024. Augmented Object Intelligence with XR-Objects. (2024). https://doi.org/10.48550/ARXIV.2404.13274
@@
- Gruenefeld et al. (2022) Uwe Gruenefeld, Jonas Auda, Florian Mathis, Stefan Schneegass, Mohamed Khamis, Jan Gugenheimer, and Sven Mayer. 2022. VRception: Rapid Prototyping of Cross-Reality Systems in Virtual Reality. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (CHI ’22). Association for Computing Machinery, New York, NY, USA, 1–15. https://doi.org/10.1145/3491102.3501821
@@
- Ipsita et al. (2021) Ananya Ipsita, Hao Li, Runlin Duan, Yuanzhi Cao, Subramanian Chidambaram, Min Liu, and Karthik Ramani. 2021. VRFromX: From Scanned Reality to Interactive Virtual Experience with Human-in-the-Loop. In Extended Abstracts of the 2021 CHI Conference on Human Factors in Computing Systems (CHI EA ’21). Association for Computing Machinery, New York, NY, USA, 1–7. https://doi.org/10.1145/3411763.3451747
@@
- Jeong et al. (2023) Yunwoo Jeong, Hyungjun Cho, Taewan Kim, and Tek-Jin Nam. 2023. AutomataStage: an AR-mediated Creativity Support Tool for Hands-on Multidisciplinary Learning. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (CHI ’23). Association for Computing Machinery, New York, NY, USA, 1–16. https://doi.org/10.1145/3544548.3581408
@@
- Kim et al. (2019) Yoonji Kim, Youngkyung Choi, Hyein Lee, Geehyuk Lee, and Andrea Bianchi. 2019. VirtualComponent: A Mixed-Reality Tool for Designing and Tuning Breadboarded Circuits. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems. ACM, Glasgow Scotland Uk, 1–13. https://doi.org/10.1145/3290605.3300407
- Lafreniere and Grossman (2018) Ben Lafreniere and Tovi Grossman. 2018. Blocks-to-CAD: A Cross-Application Bridge from Minecraft to 3D Modeling. In Proceedings of the 31st Annual ACM Symposium on User Interface Software and Technology (UIST ’18). Association for Computing Machinery, New York, NY, USA, 637–648. https://doi.org/10.1145/3242587.3242602
@@
- Laugwitz et al. (2008) Bettina Laugwitz, Theo Held, and Martin Schrepp. 2008. Construction and Evaluation of a User Experience Questionnaire. In HCI and Usability for Education and Work, Andreas Holzinger (Ed.). Springer, Berlin, Heidelberg, 63–76. https://doi.org/10.1007/978-3-540-89350-9_6
@@
- Müller et al. (2021) Leon Müller, Ken Pfeuffer, Jan Gugenheimer, Bastian Pfleging, Sarah Prange, and Florian Alt. 2021. SpatialProto: Exploring Real-World Motion Captures for Rapid Prototyping of Interactive Mixed Reality. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems (CHI ’21). Association for Computing Machinery, New York, NY, USA, 1–13. https://doi.org/10.1145/3411764.3445560
@@
- Villanueva et al. (2021) Ana M Villanueva, Ziyi Liu, Zhengzhe Zhu, Xin Du, Joey Huang, Kylie A Peppler, and Karthik Ramani. 2021. RobotAR: An Augmented Reality Compatible Teleconsulting Robotics Toolkit for Augmented Makerspace Experiences. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems (CHI ’21). Association for Computing Machinery, New York, NY, USA, 1–13. https://doi.org/10.1145/3411764.3445726
@@
- Willis et al. (2010) Karl D.D. Willis, Cheng Xu, Kuan-Ju Wu, Golan Levin, and Mark D. Gross. 2010. Interactive fabrication: new interfaces for digital fabrication. In Proceedings of the fifth international conference on Tangible, embedded, and embodied interaction (TEI ’11). Association for Computing Machinery, New York, NY, USA, 69–72. https://doi.org/10.1145/1935701.1935716
@@
- Yue et al. (2017) Ya-Ting Yue, Xiaolong Zhang, Yongliang Yang, Gang Ren, Yi-King Choi, and Wenping Wang. 2017. WireDraw: 3D Wire Sculpturing Guided with Mixed Reality. In Proceedings of the 2017 CHI Conference on Human Factors in Computing Systems (CHI ’17). Association for Computing Machinery, New York, NY, USA, 3693–3704. https://doi.org/10.1145/3025453.3025792

### v2

@@ 6.2.3. Surveys
For the standard SUS survey, we obtained a score of 90%, which shows that TinkerXR system is within the A grade (90% or above), while Tinkercad achieved a score of 85%, which places it within the B grade (from 80% to 90%) (Sauro and Lewis, 2016). As shown in Fig. 17, the TinkerXR system has superior scores across statements such as Q3, Q9, and Q10. These findings support that TinkerXR was easier to use for users in terms of coherency. The additional questions related to immersive experience show that TinkerXR helped users to have a more natural and hands-on designing procedure.
@@
The analysis of the NASA-TLX survey with a paired t-test reveals significant differences between Tinkercad and TinkerXR across multiple workload dimensions. Participants reported lower mental demand when using TinkerXR compared to Tinkercad (t = 3.857, p = 0.004), indicating that the immersive AR environment facilitated easier comprehension and interaction during design tasks. In contrast, TinkerXR exhibited higher physical demand (t = -5.658, p <0.001) and effort (t = -2.684, p = 0.025), likely due to the reliance on hand gestures and the physical interaction required with the AR system compared to the mouse-and-keyboard interface of Tinkercad. However, temporal demand, performance, and frustration levels were comparable between the two tools, indicating that the usability benefits of TinkerXR may mitigate the added physical strain.
Additionally, considering the current capabilities and limitations, users indicated that they would use TinkerXR for future projects, with an average rating of 4.4 on a 5-point Likert scale. They also recommended this tool to other novice users, giving it an average rating of 4.7 out of 5.
Result of the NASA-TLX survey.
@@ 6.3. Qualitative Feedback
Open-ended questions in our survey, such as ”What did you like most about using the tool?”, ”What challenges did you face while using the tool?”, and ”What aspects of the tool would you most like to see improved?”, and the thematic analysis of the responses allowed us to gather deeper insights into the system’s strengths and areas for improvement under five overarching themes:
Real-World Integration and Immersive Experience
Participants consistently praised the ability to design within real environments, emphasizing the value of TinkerXR’s AR capabilities in visualizing objects in situ. For instance, P1 shared, “With TinkerXR, I could see how my design would actually look in the space it was meant for.” P4 highlighted how designing anywhere — including outdoors — opened new possibilities, such as creating projects in natural settings. The immersive experience was widely appreciated, with features like depth occlusion enhancing realism. As P10 noted, “The occlusion made everything feel so much more immersive, like the objects were truly part of the room.” The engaging and enjoyable nature of the system was also mentioned, with P3 stating, “Using the system was genuinely fun and engaging.”
Intuitive and Natural Controls
The hand-based interactions and intuitive controls stood out as significant strengths. P6 described them as “human-nature,” explaining, “Locking axes was so simple, just a left-hand finger pinch and moving along the axis I wanted.” Similarly, the snap-to-grid feature was highly valued for its precision, with P2 commenting, “The snap-to-grid feature was a game-changer for precise alignments.” These intuitive mechanisms allowed participants to focus on creativity without being hindered by complex operations.
Physical Demands and Ergonomics
While the controls were intuitive, several participants expressed concerns about the physical demands of prolonged use. P1 mentioned, “The physical demand was rather tiring,” and P2 noted, “It was a little tiring to have my arm raised for every action.” These observations suggest that extended use could lead to fatigue, especially when designing complex models.
Accessibility and Future Adoption
TinkerXR was widely commended for its accessibility, even among novice users. P6 shared, “As someone completely new to AR, 3D modeling, and 3D printing, I was surprised by how easy it was to learn.” This ease of use motivated participants to envision future applications, such as designing custom objects for their homes or gardens. P4 expressed excitement about using the system for personal projects, stating, “Even in my garden, I am already thinking about designing a birdhouse and some tree decorations.” Most participants indicated they would continue using the tool and recommend it to others, with P8 emphasizing its convenience: “Starting the print directly from the headset saves materials because I don’t have to redo prints.”
Feature Expansion and Precision Tools
Participants identified areas where additional features could enhance the workflow. Several requested better tools for manipulating small objects, with P7 suggesting a “magnifying glass tool for designing small objects.” Building on the AR ruler feature, on-screen measurements were another commonly desired functionality, as P1 noted, “Length values of the edges appearing next to the objects, just like in Tinkercad, would help.” Other recommendations included importing external 3D models and freely adjusting grid sizes. P4 also proposed more direct interaction methods, such as “grabbing objects directly with our hands.”
Overall, the feedback highlights TinkerXR’s strengths in providing an immersive, intuitive, and accessible design experience, especially for novices. However, challenges like solving the problem of physical fatigue on extended usage should be solved in the future. Addressing functionalities like small object manipulation and hand-grabbing the objects would also further enhance its usability.
@@ 7. Discussion
In this section, we evaluate the effectiveness and future potential of TinkerXR. We discuss its suitability for both novice and expert users and express limitations related to handling complex 3D geometries. We also address the impact of the AR form factor on user experience and the limitations of hand gesture recognition. Additionally, we anticipate future advancements that could improve the system’s capabilities and user interaction.
@@
Our tool was designed with novices in mind, providing an intuitive interface that simplifies both the design and fabrication of models. However, we envision that TinkerXR can also be beneficial to experts. Practitioners and professional designers can use it for rapid prototyping, concept development, and interactive presentations. The ease of use and flexibility of our system can streamline workflows, allowing experts to focus on more complex design tasks without being bogged down by technical logistics. In the future, we believe that our system can be adopted by educational institutions for teaching design principles, by industries for quick mock-ups or collaborative sessions, and by hobbyists seeking to explore 3D modeling and fabrication.
@@
In Section 5, we demonstrated our tool’s usefulness through various design examples, each showing a different use case. These objects featured distinct geometries, ranging from small details to larger-scale structures. For instance, the candy bowl model in Figure 14d is made of 32 curved lines, each made of 8 cuboids, adding up to a total of 256 basic building blocks (if none of the blocks are combined) and 2048 vertices. Our system supports up to 300 basic building blocks and a total number of vertices of 20,000 in real time. However, we observed a noticeable lag when exceeding these numbers, especially in combination operations. We anticipate this limitation can be addressed with future headsets offering higher processing power. Alternatively, more powerful headsets, e.g., Varjo XR-4141414https://varjo.com/products/xr-4/, Accessed: 4.08.2024, which is connected to a desktop computer, can support compute-intensive tasks.
While the system can create a variety of complex geometries, certain tasks remain challenging due to the lack of more advanced tools such as extrusion of 3D geometries from 2D blueprints (e.g., sketched shapes, text), parametric curves, lofting, filleting, freeform design tools similar to those in Autodesk 3ds Max151515https://www.autodesk.com/products/3ds-max, Accessed: 4.08.2024, or a collaborative structure allowing multiple users to work on the same project similar to Gravity Sketch or adding physical engagement with local users and their surroundings similar to HoloBots (Ihara et al., 2023). These features were not implemented, as our system targets novices and prioritizes simplicity. While it is arguable that adding such features inherently increases complexity, our design choices focus on maintaining an intuitive user experience.
@@
We believe that TinkerXR has the potential to adapt to other digital fabrication methods such as molding (Serrano et al., 2022; Valkeneers et al., 2019), CNC machining (Tran O’Leary et al., 2023; Tran O’Leary and Peek, 2019), or laser cutting (Dogan et al., 2023a, 2021) in the future. However, we acknowledge that this adaptation goes beyond simple file format conversion. Each method introduces unique design constraints, such as draft angles for molds, toolpath limitations for CNC milling, and material thickness or 2D geometry constraints for laser cutting. These factors would require the system to integrate CAM tools to guide users in adhering to method-specific constraints.
In this context, Roumen et al. (Baudisch et al., 2019) provide inspiration for adapting 3D design for laser cutting, although it is not a general-purpose solution. Exploring ways to integrate such techniques into TinkerXR could enable users to understand fabrication-specific constraints in real time, which we consider an exciting avenue for future work.
@@
MR headsets can be bulky; however, in our user study, participants did not report major problems due to the form factor. We used the Quest 3 because it is lightweight, making it easy to use. In the future, we imagine AR technology becoming more ubiquitous while offering a more streamlined experience with the advent of AR glasses, such as Meta Orion161616https://about.meta.com/realitylabs/orion/, Accessed: 20.01.2025. Future devices are likely to be more compact, resembling ordinary eyewear, thus enhancing user comfort and accessibility (Dogan, 2024). These advancements could lead to broader adoption and more immersive systems.
@@
While we are constrained by Meta’s hand gesture APIs, which are known to improve over time171717https://www.techradar.com/computing/virtual-reality-augmented-reality/your-meta-quest-3-is-getting-a-hand-tracking-upgrade-that-could-unlock-foot-tracking, Accessed: 5.08.2024, there are still issues such as improper functionality in dimly lit areas or limited precision of tracking compared to controllers or the system being exclusive to Meta headsets.
@@ 8. Conclusion
In this paper, we presented TinkerXR, a new 3D modeling and fabrication tool through AR, designed for novices. Our approach lets users design artifacts with a simple primitive geometry manipulation CSG design logic, with intuitive hand gesture control schemes, on the intended space of the artifact, and with seamless integration with the 3D printing process. We showed the various features present in TinkerXR from selecting the workspace to design and fabrication. Using these features, we presented a wide variety of creations for different scenarios, each varying in complexity and practical applications. A user study revealed that the system was effective for novices, and more immersive than Tinkercad, while offering a comparable level of functionality and simplicity. We hope that our project and its open-source codebase will stimulate further exploration within communities engaged in HCI, design, and fabrication.
@@ Acknowledgements.
- Bourdot et al. (2010) P. Bourdot, T. Convard, F. Picon, M. Ammi, D. Touraine, and J.-M. Vézien. 2010. VR–CAD integration: Multimodal immersive interaction and advanced haptic paradigms for implicit edition of CAD models. Computer-Aided Design 42, 5 (May 2010), 445–461. https://doi.org/10.1016/j.cad.2008.10.014
@@
- Cordeiro et al. (2019) Eduardo Cordeiro, Franca Giannini, and Marina Monti. 2019. A Survey of Immersive Systems for Shape Manipulation. Computer-Aided Design and Applications 16, 6 (March 2019), 1146–1157. https://doi.org/10.14733/cadaps.2019.1146-1157
- Dogan (2024) Mustafa Doga Dogan. 2024. Ubiquitous Metadata: Design and Fabrication of Embedded Markers for Real-World Object Identification and Interaction. https://doi.org/10.48550/arXiv.2407.11748 arXiv:2407.11748 [cs].
- Dogan et al. (2022a) Mustafa Doga Dogan, Patrick Baudisch, Hrvoje Benko, Michael Nebeling, Huaishu Peng, Valkyrie Savage, and Stefanie Mueller. 2022a. Fabricate It or Render It? Digital Fabrication vs. Virtual Reality for Creating Objects Instantly. In Extended Abstracts of the 2022 CHI Conference on Human Factors in Computing Systems. Association for Computing Machinery, 5. https://doi.org/10.1145/3491101.3516510
- Dogan et al. (2023a) Mustafa Doga Dogan, Vivian Hsinyueh Chan, Richard Qi, Grace Tang, Thijs Roumen, and Stefanie Mueller. 2023a. StructCode: Leveraging Fabrication Artifacts to Store Data in Laser-Cut Objects. In Proceedings of the 8th ACM Symposium on Computational Fabrication (SCF ’23). Association for Computing Machinery, New York, NY, USA, 1–13. https://doi.org/10.1145/3623263.3623353
- Dogan et al. (2021) Mustafa Doga Dogan, Steven Vidal Acevedo Colon, Varnika Sinha, Kaan Akşit, and Stefanie Mueller. 2021. SensiCut: Material-Aware Laser Cutting Using Speckle Sensing and Deep Learning. In Proceedings of the 34th Annual ACM Symposium on User Interface Software and Technology. ACM, Virtual Event USA, 15. https://doi.org/10.1145/3472749.3474733
- Dogan et al. (2023b) Mustafa Doga Dogan, Raul Garcia-Martin, Patrick William Haertel, Jamison John O’Keefe, Ahmad Taka, Akarsh Aurora, Raul Sanchez-Reillo, and Stefanie Mueller. 2023b. BrightMarker: 3D Printed Fluorescent Markers for Object Tracking. In Proceedings of the 36th Annual ACM Symposium on User Interface Software and Technology (UIST ’23). Association for Computing Machinery, New York, NY, USA, 1–13. https://doi.org/10.1145/3586183.3606758
- Dogan et al. (2024) Mustafa Doga Dogan, Eric J Gonzalez, Karan Ahuja, Ruofei Du, Andrea Colaço, Johnny Lee, Mar Gonzalez-Franco, and David Kim. 2024. Augmented Object Intelligence with XR-Objects. In Proceedings of the 37th Annual ACM Symposium on User Interface Software and Technology (UIST ’24). Association for Computing Machinery, New York, NY, USA, 1–15. https://doi.org/10.1145/3654777.3676379
- Dogan et al. (2023c) Mustafa Doga Dogan, Alexa F. Siu, Jennifer Healey, Curtis Wigington, Chang Xiao, and Tong Sun. 2023c. StandARone: Infrared-Watermarked Documents as Portable Containers of AR Interaction and Personalization. In Extended Abstracts of the 2023 CHI Conference on Human Factors in Computing Systems (CHI EA ’23). Association for Computing Machinery, New York, NY, USA, 1–7. https://doi.org/10.1145/3544549.3585905
- Dogan et al. (2022b) Mustafa Doga Dogan, Ahmad Taka, Michael Lu, Yunyi Zhu, Akshat Kumar, Aakar Gupta, and Stefanie Mueller. 2022b. InfraredTags: Embedding Invisible AR Markers and Barcodes Using Low-Cost, Infrared-Based 3D Printing and Imaging Tools. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems. Association for Computing Machinery, New Orleans LA USA, 9. https://doi.org/10.1145/3491102.3501951
- Dogangun et al. (2024) Fatih Dogangun, Serdar Bahar, Yigit Yildirim, Bora Toprak Temir, Emre Ugur, and Mustafa Doga Dogan. 2024. RAMPA: Robotic Augmented Reality for Machine Programming and Automation. https://doi.org/10.48550/arXiv.2410.13412 arXiv:2410.13412 [cs].
@@
- Feeman et al. (2018) Seth M. Feeman, Landon B. Wright, and John L. Salmon. 2018. Exploration and evaluation of CAD modeling in virtual reality. Computer-Aided Design and Applications 15, 6 (Nov. 2018), 892–904. https://doi.org/10.1080/16864360.2018.1462570
@@
- Hofmann et al. (2018) Megan Hofmann, Gabriella Hann, Scott E. Hudson, and Jennifer Mankoff. 2018. Greater than the Sum of its PARTs: Expressing and Reusing Design Intent in 3D Models. In Proceedings of the 2018 CHI Conference on Human Factors in Computing Systems. ACM, Montreal QC Canada, 1–12. https://doi.org/10.1145/3173574.3173875
@@
- Iyer et al. (2025) Ramya Iyer, Mustafa Doga Dogan, Maria Larsson, and Takeo Igarashi. 2025. XR-penter: Material-Aware and In Situ Design of Scrap Wood Assemblies. In Proceedings of the 2025 CHI Conference on Human Factors in Computing Systems (CHI ’25). Association for Computing Machinery, New York, NY, USA.
@@
- Jasche and Ludwig (2020) Florian Jasche and Thomas Ludwig. 2020. PrintARface: Supporting the Exploration of Cyber-Physical Systems through Augmented Reality. In Proceedings of the 11th Nordic Conference on Human-Computer Interaction: Shaping Experiences, Shaping Society. ACM, Tallinn Estonia, 1–12. https://doi.org/10.1145/3419249.3420162
@@
- Jiang et al. (2021) Ying Jiang, Congyi Zhang, Hongbo Fu, Alberto Cannavò, Fabrizio Lamberti, Henry Y K Lau, and Wenping Wang. 2021. HandPainter - 3D Sketching in VR with Hand-based Physical Proxy. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems (CHI ’21). Association for Computing Machinery, New York, NY, USA, 1–13. https://doi.org/10.1145/3411764.3445302
@@
- Lau et al. (2012) Manfred Lau, Masaki Hirose, Akira Ohgawara, Jun Mitani, and Takeo Igarashi. 2012. Situated modeling: a shape-stamping interface with tangible primitives. In Proceedings of the Sixth International Conference on Tangible, Embedded and Embodied Interaction. ACM, Kingston Ontario Canada, 275–282. https://doi.org/10.1145/2148131.2148190
@@
- Nuernberger et al. (2016) Benjamin Nuernberger, Eyal Ofek, Hrvoje Benko, and Andrew D. Wilson. 2016. SnapToReality: Aligning Augmented Reality to the Real World. In Proceedings of the 2016 CHI Conference on Human Factors in Computing Systems. ACM, San Jose California USA, 1233–1244. https://doi.org/10.1145/2858036.2858250
@@
- Rosales et al. (2021) Enrique Rosales, Chrystiano Araújo, Jafet Rodriguez, Nicholas Vining, Dongwook Yoon, and Alla Sheffer. 2021. AdaptiBrush: adaptive general and predictable VR ribbon brush. ACM Trans. Graph. 40, 6 (Dec. 2021), 247:1–247:15. https://doi.org/10.1145/3478513.3480511
@@
- Sauro and Lewis (2016) Jeff Sauro and James R. Lewis. 2016. Quantifying the user experience: Practical statistics for user research. Morgan Kaufmann.
- Serrano et al. (2022) Marcos Serrano, Jolee Finch, Pourang Irani, Andrés Lucero, and Anne Roudaut. 2022. Mold-It: Understanding how Physical Shapes affect Interaction with Handheld Freeform Devices. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (CHI ’22). Association for Computing Machinery, New York, NY, USA, 1–14. https://doi.org/10.1145/3491102.3502022
@@
- Tran O’Leary et al. (2023) Jasper Tran O’Leary, Gabrielle Benabdallah, and Nadya Peek. 2023. Imprimer: Computational Notebooks for CNC Milling. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (CHI ’23). Association for Computing Machinery, New York, NY, USA, 1–15. https://doi.org/10.1145/3544548.3581334
- Tran O’Leary and Peek (2019) Jasper Tran O’Leary and Nadya Peek. 2019. Machine-o-Matic: A Programming Environment for Prototyping Digital Fabrication Workflows. In Adjunct Proceedings of the 32nd Annual ACM Symposium on User Interface Software and Technology (UIST ’19 Adjunct). Association for Computing Machinery, New York, NY, USA, 134–136. https://doi.org/10.1145/3332167.3356897
- Valkeneers et al. (2019) Tom Valkeneers, Danny Leen, Daniel Ashbrook, and Raf Ramakers. 2019. StackMold: Rapid Prototyping of Functional Multi-Material Objects with Selective Levels of Surface Details. In Proceedings of the 32nd Annual ACM Symposium on User Interface Software and Technology (UIST ’19). Association for Computing Machinery, New York, NY, USA, 687–699. https://doi.org/10.1145/3332165.3347915